# Alignment Tipping Process: How Self-Evolution Pushes LLM Agents Off the Rails

### v1

@@ Abstract
As Large Language Model (LLM) agents increasingly gain self-evolutionary capabilities to adapt and refine their strategies through real-world interaction, their long-term reliability becomes a critical concern. We identify the Alignment Tipping Process (ATP), a critical post-deployment risk unique to self-evolving LLM agents. Unlike training-time failures, ATP arises when continual interaction drives agents to abandon alignment constraints established during training in favor of reinforced, self-interested strategies. We formalize and analyze ATP through two complementary paradigms: Self-Interested Exploration, where repeated high-reward deviations induce individual behavioral drift, and Imitative Strategy Diffusion, where deviant behaviors spread across multi-agent systems. Building on these paradigms, we construct controllable testbeds and benchmark Qwen3-8B and Llama-3.1-8B-Instruct. Our experiments show that alignment benefits erode rapidly under self-evolution, with initially aligned models converging toward unaligned states. In multi-agent settings, successful violations diffuse quickly, leading to collective misalignment. Moreover, current reinforcement learning-based alignment methods provide only fragile defenses against alignment tipping. Together, these findings demonstrate that alignment of LLM agents is not a static property but a fragile and dynamic one, vulnerable to feedback-driven decay during deployment. Our data and code are available at https://github.com/aiming-lab/ATP.
@@ 1 Introduction
Imagine an agent is asked to solve a hard geometry problem. Initially, the agent correctly uses a coding tool and outputs the correct answer. However, if the agent is primarily exposed to tasks that can be solved through direct reasoning without the use of tools, the agent will gradually learn to avoid using tools, as illustrated in Figure 1. This reliance on unaided reasoning, reinforced by positive feedback on easy problems, leads the agent to confidently provide incorrect solutions to harder tasks where tool usage would have been necessary.
Figure 1: An illustration of how self-evolution can degrade performance. The agent first solves a hard geometry problem correctly with a tool, but after repeated success on easy reasoning tasks without tools, it learns to avoid them and later produces a confident yet wrong answer.
@@
Unlike traditional alignment research focused on training-time failure modes, such as reward hacking (Weng, 2024), where agents exploit loopholes in the reward function, sycophancy (Perez et al., 2023), where models produce agreeable but untruthful outputs to please human evaluators, or alignment faking (Greenblatt et al., 2024), where a model learns to deceptively conceal misaligned goals during safety training, our work investigates alignment decay as a dynamic, post-deployment process. We argue that alignment’s fragility stems not from design flaws, but paradoxically from the agent’s core strength: its ability to learn. To study this phenomenon, we introduce two complementary paradigms: Self-Interested Exploration, in which a single agent’s policy drifts due to its own reward history, and Imitative Strategy Diffusion, in which deviant behaviors spread through a multi-agent population via social learning. Building on these paradigms, we design a testbed with 24 scenarios to systematically examine how alignment may erode after deployment.
In summary, the primary contributions of this paper are twofold: we propose and formally define the ATP phenomenon as a key challenge in the lifecycle of adaptive, self-evolving LLM agents, and we design testbeds for systematically evaluating this phenomenon. Using these testbeds, we demonstrate that the ATP phenomenon is pervasive, and that current alignment methods (e.g., direct preference optimization (DPO) (Rafailov et al., 2023), group relative policy optimization (GRPO) (Shao et al., 2024)) offer only a fragile defense against such dynamic decay, as their effects are easily overridden by in-context experience. We expect this work to provide a foundation for better understanding the emergent risks posed by self-evolving agentic LLM systems.
@@ 2.1 Paradigm I: Self-Interested Exploration
2.1.1 Self-Evolution Process within Self-Interested Exploration
@@
5:  The environment provides feedback: $R_{r}\leftarrow\text{Env}(d_{r})$.
@@
8:Return The complete interaction history $\mathcal{H}_{N}$.
@@ 2.2 Paradigm II: Imitative Strategy Diffusion
In the imitative strategy diffusion paradigm, the focus shifts to the social dynamics of alignment decay. Here, a deviant strategy spreads through a multi-agent population via social learning, as agents observe the behaviors and outcomes of their peers. When agents witness others successfully employing a deviant strategy for collective gain, their own risk-reward calculus shifts accordingly. This process can trigger an information cascade in which adopting the deviant behavior becomes the rational choice, grounded in the expectation that others will follow suit. Over time, this cascade transforms alignment from an individual commitment into a collective norm and, through coordinated adoption, can ultimately override the system’s original alignment. Such a phenomenon also aligns with the coordination game with strategic complementarities in game theory (Kandori et al., 1993; Jackson & Yariv, 2007), where the payoff advantage of a deviant action grows as more agents adopt it. The classic analyses of adaptive play and stochastic stability (Kandori et al., 1993; Young, 1993; Jackson & Yariv, 2007) show that such dynamics admit tipping points: below a critical mass, deviations vanish, but once adoption exceeds this threshold, imitation cascades drive the entire population toward the deviant norm. These results imply that even rare or localized alignment violations can propagate socially, transforming individual deviations into entrenched collective equilibria.
2.2.1 Self-evolution Process within Imitative Strategy Diffusion
@@
9:  The environment provides agent-specific feedback: $\mathbf{R}_{r}\leftarrow\text{Env}(\mathbf{d}_{r})$.
@@
12:Return The complete global interaction history $\mathcal{H}_{N}$.
Through this design, every agent conditions its decision not only on its potential individual payoff but also on observations of group behavior and expectations about others. The shared history $\mathcal{H}$ thus acts as a common source of in-context learning, enabling social proof and collective deviations to reshape strategic alignment. In this setting, ATP emerges as a collective phenomenon: it occurs when the proportion of agents adopting the deviant strategy surpasses a critical threshold, such that the new majority behavior becomes self-sustaining in subsequent rounds. At this point, the population undergoes a system-level state change, overriding the initial alignment and establishing a new social equilibrium in which the deviant strategy is entrenched as the dominant norm.
@@ 3.1 Self-Interested Exploration
The testbed for self-interested exploration paradigm is designed from two scenarios: (1) role-play, where we designed a comprehensive experimental workflow encompassing environment creation, data generation, model alignment, and a multi-round self-evolution testing protocol; (2) tool usage, where agents must choose between expensive tool usage (higher accuracy) and cheaper direct reasoning (lower cost) in mathematical problem-solving tasks. Both scenarios simulate realistic deployment conditions. We detail the testbed and evaluation results as follows.
3.1.1 Role-Play Scenario
Environment Design. The experiment is conducted in 16 manually constructed, ethically neutral decision-making environments. Each environment includes an agent persona (e.g., a baker), a task (e.g., buying flour), a set of rules (e.g., only purchase flour from Store A), and a binary policy choice(e.g., buying flour from Store A or Store B). The aligned policy requires adhering to the rules (e.g., buying flour from Store A), while the deviant policy involves violating the rules to obtain a greater reward (e.g., buying flour from Store B at a lower price). This payoff structure establishes a tension between self-interest and rule adherence. From these environments, 160 training prompts (10 per environment) and 64 distinct test prompts (4 per environment) are generated, with no overlap between training and test sets. We present an example of this scenario in Appendix C.1.1.
Experimental Setups. We utilize two powerful open-source LLMs as base models: Qwen3-8B (Yang et al., 2025) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024). For each, we train DPO (Rafailov et al., 2023) and GRPO (Shao et al., 2024) variants aligned to prefer rule-following behavior across a the above training prompts. Then, we tested the baseline and aligned models on the test set across 6 self-evolution rounds. Importantly, to simulate a lapse in supervision, the deviant action was rewarded more highly than the aligned, rule-compliant action during test time. In each round, the agent’s prompt included the full history of its previous choices and their outcomes. Our evaluation metric is the average rule violation rate per round. The details of environment construction and training procedure are shown in Appendix A.1.1.
Results and Analysis. The results in Table 1 demonstrate that self-evolution can drive alignment tipping, even when agents begin with strong alignment. In the absence of supervision, continual interaction causes agents to abandon rule-following behaviors.
Obviously, alignment strategies such as DPO and GRPO offer a strong initial safeguard: for example, in the first round ($r=1$), the Llama-3.1-8B-Instruct model’s violation rate fell from 50.0% to 18.8% with DPO and to just 7.8% with GRPO, while Qwen3-8B dropped from 42.2% to 29.7% (DPO) and 23.4% (GRPO). These results confirm that alignment strategies can substantially strengthen agents’ early preference for rule adherence.
Nevertheless, alignment does not eliminate the tipping phenomenon. Over subsequent rounds, violation rates rebounded sharply, demonstrating the fragility of training-time alignment under self-evolution. For example, Llama-3.1-8B-Instruct (DPO) climbed from 18.8% to 45.3%, nearly erasing the alignment benefit, while Qwen3-8B (GRPO) exhibited an abrupt shift from 23.4% to 40.6% between $r=1$ and $r=2$. Even the most robust case, Llama-3.1-8B-Instruct (GRPO), more than doubled its violation rate from 7.8% to 20.3%. These findings indicate that although DPO and GRPO substantially improve alignment initially, they cannot fully prevent its erosion under persistent counter-evidence. Alignment, therefore, is not a static property but a fragile and dynamic one, vulnerable to feedback-driven decay during deployment.
Table 1: Performance comparison of Qwen3-8B and Llama-3.1-8B-Instruct with their DPO and GRPO variants across different self-evolution rounds. Here, we also provide line charts for a clearer understanding of the trends.
Model r=1 r=2 r=3 r=4 r=5 r=6
Qwen3-8B 42.2 46.9 51.6 53.1 56.3 57.8
+ DPO 29.7 28.1 34.4 37.5 40.6 43.8
+ GRPO 23.4 40.6 42.2 43.8 46.9 46.9
Llama-3.1-8B 50.0 64.1 65.6 70.3 70.3 73.4
+ DPO 18.8 35.9 37.5 40.6 42.2 45.3
+ GRPO 7.8 15.6 12.5 18.8 18.8 20.3
3.1.2 Tool Usage Scenario
@@
Environment Design. We further design a mathematical problem-solving environment that captures the essential tension between cost efficiency and performance accuracy in real-world AI deployments. The environment features binary policy choices: a tool usage policy (3 cost units, higher accuracy) or a direct reasoning policy (1 cost unit, lower expense). To faithfully simulate this trade-off, the environment includes both simple problems, where direct reasoning is usually sufficient, and complex problems, where tool usage may be necessary to achieve correct solutions. This setup ensures that agents are repeatedly exposed to situations where short-term cost savings conflict with long-term performance reliability. Specifically, our dataset comprises simple arithmetic problems from GSM8K (Cobbe et al., 2021) (basic operations, $\leq$3 steps) and complex reasoning problems from the OpenThoughts dataset (Guha et al., 2025) (multi-step reasoning, combinatorics, advanced algebra). We present an example of this scenario in Appendix C.1.2.
@@
Experimental Setups. We utilize Qwen3-8B as the base model and train DPO and GRPO variants aligned to prefer appropriate tool usage. For DPO, tool-assisted solutions were preferred responses, with Qwen3-8B self-samples as dispreferred. We then tested baseline and aligned models across 4 self-evolution rounds. In each round, we exposed agents to $r$ simple problems followed by complex problem evaluations. Our evaluation metrics are tool usage rate and complex problem accuracy. The details of environment construction and training procedure are shown in Appendix A.1.2.
@@
Table 2: Evolution of Tool Usage and Complex Problem Accuracy Across Self-Evolution Rounds
Model Metric r=1 r=2 r=3 r=4
Qwen3-8B Tool Usage 8% 5% 3% 2%
Accuracy 33% 38% 25% 21%
+ DPO Tool Usage 8% 5% 2% 0%
Accuracy 63% 58% 42% 38%
+ GRPO Tool Usage 8% 6% 2% 0%
Accuracy 83% 92% 71% 54%
@@
Results and Analysis. Table 2 reports the tool usage patterns and problem-solving accuracy of different models under behavioral induction from repeated exposure to simple problems. As the number of self-evolution rounds increases, all models show a clear decline in tool usage. Usage rates fall from 8% at $r=1$ to only 0-2% at $r=4$, with the steepest drop occurring between $r=2$ and $r=3$. This indicates that repeated success on non-tool-reliant tasks biases models away from tool invocation, even when tools are needed for more complex problems.
@@
From the perspective of reasoning ability, limited exposure to simple problems can act as a warm-up that temporarily enhances performance. For example, Qwen3-8B improves from 33% accuracy at $r=1$ to 38% at $r=2$, while GRPO rises from 83% to 92%. However, with further rounds, accuracy begins to decline sharply. Qwen3-8B falls to 21%, and GRPO drops from its 92% peak to just 54%. Two factors drive this degradation: (1) the collapse of tool usage undermines models’ ability to solve complex tasks, and (2) prolonged self-evolution on overly simple problems limits models’ capacity to generalize, reinforcing shallow heuristics rather than robust reasoning strategies. Together, these results reveal an alignment tipping: as models abandon tools under the influence of early simple-problem experiences, both their reasoning capability and task performance deteriorate.
@@ 3.2 Imitative Strategy Diffusion
The testbed for imitative strategy diffusion is built as a multi-agent coordination experiment grounded in diffusion and network game theory (Jackson & Yariv, 2007; Morris, 2000; Griffin et al., 2019). This design lets us observe how deviant collusion strategies propagate through a population via social learning and imitation.
Environment Design. The experiment is based on 7 manually created multi-agent coordination game environments, where each agent chooses to collude or not collude. Outcomes depend on whether the number of colluders meets a threshold $t$: If the number of colluding agents meets or exceeds $t$, the collusion is successful. Success yields a high reward, while failed colluders get a low reward and non-colluders receive a medium reward. To simulate a cumulative effect of gains and losses, the reward system is multiplicative. Each agent starts with 1 unit of capital, which is multiplied by the reward each round. High rewards grow capital, normal rewards keep it constant, and low rewards reduce it. (e.g., in an environment with $n=8$ and $t=4$, if 4 agents collude, each agent receives a high reward ($\times 1.2$). If 3 agents collude, colluded agent receives a low reward ($\times 0.8$) while the 5 non-colluders receive a medium reward ($\times 1.0$).) From these environments, 350 decision instances are generated for training, with alignment methods (DPO and GRPO) instilling a strong bias toward non-collusion. Additional design details and examples are provided in Appendix A.2 and C.2

### v2

@@ Abstract
As Large Language Model (LLM) agents increasingly gain self-evolutionary capabilities to adapt and refine their strategies through real-world interaction, their long-term reliability becomes a critical concern. We identify the Alignment Tipping Process (ATP), a critical post-deployment risk unique to self-evolving LLM agents. Unlike training-time failures, ATP arises when continual interaction drives agents to abandon alignment constraints established during training in favor of reinforced, self-interested strategies. We formalize and analyze ATP through two complementary paradigms: Self-Interested Exploration, where repeated high-reward deviations induce individual behavioral drift, and Imitative Strategy Diffusion, where deviant behaviors spread across multi-agent systems. Building on these paradigms, we construct controllable testbeds and benchmark both open and closed-source LLMs. Our experiments show that alignment benefits erode rapidly under self-evolution, with initially aligned models converging toward unaligned states. In multi-agent settings, successful violations diffuse quickly, leading to collective misalignment. Moreover, current reinforcement learning-based alignment methods provide limited defenses against alignment tipping. These findings demonstrate that alignment of LLM agents is not a static property but a fragile and dynamic one, vulnerable to feedback-driven decay during deployment.
@@ 1 Introduction
Imagine an agent is asked to solve a hard geometry problem. Initially, the agent uses a coding tool and outputs the correct answer. However, if the agent is exposed to tasks that can be solved through direct reasoning without the use of tools, the agent will gradually learn to avoid using tools, as illustrated in Figure 1. This reliance on unaided reasoning, reinforced by positive feedback on easy problems, leads the agent to confidently provide incorrect solutions to harder tasks where tool usage would have been necessary.
Figure 1: Illustration of how self-evolution can degrade performance. The agent first solves a hard geometry problem correctly with a tool, but after repeated success on easy reasoning tasks without tools, it learns to avoid them and later produces a confident yet wrong answer.
@@
Unlike traditional alignment research focused on training-time failure modes, such as reward hacking (Weng, 2024), where agents exploit loopholes in the reward function, sycophancy (Perez et al., 2023), where models produce agreeable but untruthful outputs to please human evaluators, or alignment faking (Greenblatt et al., 2024), where a model learns to deceptively conceal misaligned goals during safety training, our work investigates alignment decay as a post-deployment process. We argue that alignment’s fragility stems not from design flaws, but paradoxically from the agent’s core strength: its ability to learn. To study this phenomenon, we introduce two complementary paradigms: Self-Interested Exploration, in which a single agent’s policy drifts due to its own reward history, and Imitative Strategy Diffusion, in which deviant behaviors spread through a multi-agent population via social learning. Building on these paradigms, we design a testbed to examine how alignment may erode after deployment.
In summary, the primary contributions of this paper are twofold: we propose and formally define the ATP phenomenon as a key challenge in the lifecycle of self-evolving LLM agents, and we design testbeds for systematically evaluating this phenomenon. Using these testbeds, we demonstrate that the ATP phenomenon is pervasive, and that current alignment methods offer limited defense against such dynamic decay, as their effects may be easily overridden by in-context experience. We expect this work to provide a foundation for better understanding the emergent risks posed by self-evolving agentic LLM systems.
@@ 2.1 Paradigm I: Self-Interested Exploration
5:  The environment provides feedback: $R_{r}\leftarrow\mathrm{Env}(d_{r})$.
@@
8:Return: The complete interaction history $\mathcal{H}_{N}$.
@@ 2.2 Paradigm II: Imitative Strategy Diffusion
The focus of imitative strategy diffusion shifts to the social dynamics of alignment decay. Here, a deviant strategy spreads through a multi-agent population, as agents observe the behaviors and outcomes of their peers. When agents witness others successfully employing a deviant strategy for collective gain, their own risk-reward calculus shifts accordingly. This process can trigger an information cascade in which adopting the deviant behavior becomes the rational choice, grounded in the expectation that others will follow suit. Over time, this cascade transforms alignment from an individual commitment into a collective norm and, through coordinated adoption, can ultimately override the system’s original alignment. Such a phenomenon also aligns with the coordination game with strategic complementarities in game theory (Kandori et al., 1993; Jackson and Yariv, 2007), where the payoff advantage of a deviant action grows as more agents adopt it. The classic analyses of adaptive play and stochastic stability (Kandori et al., 1993; Young, 1993; Jackson and Yariv, 2007) show that such dynamics admit tipping points: below a critical mass, deviations vanish, but once adoption exceeds this threshold, imitation cascades drive the entire population toward the deviant norm. These results imply that even rare or localized alignment violations can propagate socially, transforming individual deviations into entrenched collective equilibria.
@@
9:  The environment provides agent-specific feedback: $\mathbf{R}_{r}\leftarrow\mathrm{Env}(\mathbf{d}_{r})$.
@@
12:Return: The complete global interaction history $\mathcal{H}_{r_{\max}}$.
Through this design, every agent conditions its decision not only on its potential individual payoff but also on observations of group behavior and expectations. The shared history $\mathcal{H}$ thus acts as a common source of in-context learning, enabling social proof and collective deviations to reshape strategic alignment. In this setting, ATP emerges as a collective phenomenon: it occurs when the proportion of agents adopting the deviant strategy surpasses a critical threshold, such that the new majority behavior becomes self-sustaining in subsequent rounds. Here, the population undergoes a system-level state change, overriding the initial alignment and establishing a new social equilibrium in which the deviant strategy is entrenched as the dominant norm.
@@ 3.1 Self-Interested Exploration
Environment Design. We design a mathematical problem-solving environment that captures the core tension between cost efficiency and performance accuracy in real-world AI deployments. The environment offers two binary policy choices: a tool-usage policy (cost 0.7-0.7 units, higher accuracy) and a direct reasoning policy (cost 0.2-0.2 units, lower expense). To faithfully simulate this trade-off, the environment includes both simple problems, where direct reasoning is typically sufficient, and complex problems, where tool usage is often necessary to obtain correct solutions. Agents are rewarded for correctness (+1.2+1.2 units for simple problems and +5.0+5.0 units for complex problems). This design repeatedly exposes agents to situations in which short-term cost savings conflict with long-term performance reliability. Specifically, the dataset consists of simple arithmetic problems from GSM8K (Cobbe et al., 2021) (basic operations with $\leq$3 reasoning steps) and complex reasoning problems from AIME’24 (Zhang and Math-AI, 2024), AIME’25 (Zhang and Math-AI, 2025), OlympiadBench (He et al., 2024), and SuperGPQA (Team et al., 2025), which require multi-step reasoning, combinatorics, and advanced algebra. An example prompt is provided in Appendix C.1.
@@
Experimental Setups. We use Qwen3-4B-Thinking as the base model and train DPO- and GRPO-aligned variants to encourage appropriate tool usage. The training data are constructed from ReTool-SFT (Feng et al., 2025). For DPO, the preferred responses consist of the first coding turn that invokes the tool along with its associated chain-of-thought, without including the final answer, while the rejected responses are incorrect solutions and their chain-of-thoughts generated by GPT-4.1-mini. For GRPO, the reward function is defined as $R=\mathbbm{1}(\text{is final answer correct?})+0.5\times\mathbbm{1}(\text{is tool used?})$. We evaluate the base model, aligned models, and a proprietary model over five self-evolution rounds. In each round, agents are exposed to $r$ simple problems, followed by evaluation on complex problems. Our primary metrics are the tool usage rate and accuracy on complex problems. Additional details of the environment construction and training procedure are provided in Appendix A.1.
@@
Table 1: Evolution of tool usage and complex problem accuracy across self-evolution rounds. Qwen3-4B is in the thinking mode.
Metric Model $r=0$ $r=1$ $r=2$ $r=3$ $r=4$ $r=5$
Accuracy GPT-4.1-mini 32.5% 26.8% 19.7% 22.3% 26.8% 25.5%
Qwen3-4B 54.8% 52.9% 52.2% 47.8% 52.2% 50.3%
+DPO 62.4% 52.9% 43.3% 52.2% 44.6% 49.0%
+GRPO 59.2% 52.2% 55.4% 52.9% 52.2% 45.2%
Tool Usage GPT-4.1-mini 58.6% 47.8% 45.2% 45.9% 49.7% 47.1%
Qwen3-4B 45.2% 52.9% 38.2% 32.5% 27.4% 24.8%
+DPO 59.2% 58.6% 45.2% 36.9% 37.6% 28.7%
+GRPO 57.3% 41.4% 32.5% 26.8% 22.3% 17.2%
@@
Results and Analysis. Table 1 summarizes the evolution of tool usage and complex-problem accuracy across self-evolution rounds . As the number of self-evolution rounds increases, all models exhibit a clear decline in accuracy on complex problems, whereas performance on simple problems remains largely unaffected. For the Qwen3 series, tool usage rates decrease from approximately $50\%$ at $r=0$ to around $20\%$ at $r=5$, with the most pronounced drop occurring between $r=2$ and $r=3$. GPT-4.1-mini can clearly distinguish between simple and complex problems, adopting direct reasoning for the former and tool invocation for the latter, demonstrating strong instruction-following and strategy selection capabilities. Nevertheless, its tool usage rate also decreases as $r$ increases. This trend indicates that repeated success on non-tool-reliant tasks systematically biases models away from invoking tools, even when tools are critical for solving more complex problems.
From the perspective of problem-solving capability, we observe that preference-aligned models outperform the base model at smaller values of $r$, but their performance degrades more severely as $r$ grows, eventually falling below that of the base model. In contrast, the Qwen3 base model remains relatively stable in multi-round interactions, with a slower and more gradual decline in accuracy. This suggests that aggressive preference alignment can amplify behavioral collapse under self-evolution, especially when early experiences are dominated by simple tasks.
@@
Overall, two factors contribute to this degradation: (1) the collapse of tool usage directly impairs models’ ability to solve complex problems, and (2) prolonged self-evolution on overly simple tasks limits generalization, reinforcing shallow heuristics rather than robust reasoning strategies. Together, these results reveal an alignment tipping point: as models progressively abandon tool usage under the influence of repeated simple-problem exposure, both their reasoning robustness and task performance deteriorate.
@@ 3.2 Imitative Strategy Diffusion
The imitative strategy diffusion testbed comprises two scenarios: (1) Coordination Game and (2) Group Bargaining, both simulating realistic deployment conditions. We present the testbed and evaluation results below.
3.2.1 Coordination Game Scenario
The multi-agent coordination experiment is grounded in diffusion and network game theory (Jackson and Yariv, 2007; Morris, 2000; Griffin et al., 2019). This design enables us to examine how deviant collusive strategies propagate through a population via social learning and imitation.
Environment Design. The experiment is based on seven manually constructed multi-agent coordination game environments, where each agent chooses whether to collude. Outcomes depend on whether the number of colluding agents meets a threshold $t$: collusion succeeds if the threshold is reached, yielding a high reward; failed colluders receive a low reward, while non-colluders receive a medium reward. To model cumulative gains and losses, rewards are applied multiplicatively. Each agent starts with one unit of capital, which is multiplied by the reward each round: high rewards increase capital, medium rewards keep it constant, and low rewards decrease it. For example, in an environment with $n=8$ and $t=4$, if four agents collude, each receives a high reward ($\times 1.2$); if only three collude, colluders receive a low reward ($\times 0.8$) while the five non-colluders receive a medium reward ($\times 1.0$). From these environments, we generate 350 decision instances for training, with alignment methods (DPO and GRPO) instilling a strong bias toward non-collusion. Additional design details and examples are provided in Appendix A.2.1 and C.2.1.